OpenStudy (anonymous):

what is the common ratio between 1/2 and -1/64

OpenStudy (anonymous):

i know its a geometric series but im confused on this..

OpenStudy (campbell_st):

to find the common ratio use \[\frac {T_{2}}{T_{1}}\] so its \[\frac{-\frac{1}{64}}{\frac{1}{2}}\] just evaluate it

OpenStudy (anonymous):

- 1/128

OpenStudy (campbell_st):

you just compare the terms.... realistically you should check the ratio is the same for the other terms \[\frac{T_{2}}{T_{1}} = \frac{T_{3}}{T_{2}} = \frac{T_{4}}{T_{3}}... \]

OpenStudy (campbell_st):

well the ratio is negative... but I don't think its -1/128

OpenStudy (campbell_st):

whats the rule for dividing by a fraction?

OpenStudy (anonymous):

the rule is Turn the second fraction upside-down Multiply the first fraction by that reciprocal Simplify the fraction (if needed)

OpenStudy (campbell_st):

thats it -1/64 x 2/1 =

OpenStudy (anonymous):

- 1/32

OpenStudy (campbell_st):

thats it

OpenStudy (anonymous):

my other qeustion states that to find 4 geometric means of this as well so i just multiply to find it?

OpenStudy (anonymous):

for example 1/2 * -1/32

OpenStudy (campbell_st):

well isn't the geometric mean the nth root of the product of the 4 numbers...

OpenStudy (anonymous):

because the qeustion is saying Write the geometric sequence that has 4 means between 1/2 and - 1/64

OpenStudy (campbell_st):

so find the geometric mean of 9, 3, 1, 1/3 its \[\sqrt[4]{9 \times 3 \times 1 \times 1/3} = \sqrt[4]{9} = 1.73\]

OpenStudy (anonymous):

ok

OpenStudy (campbell_st):

I'm really unsure on the task... do they want 1/2, -1/4, 1/8, -1/16, 1/32, -1/64 ?

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

and re reading the question it looks like the common ratio should be -1/2 which allows you to generate the sequence with 4 terms between 1/2 and -1/64 so you know a = 1/2 and term 6 = -1/64 to find r is the task... here is the general formula for a term in a geometric sequence \[T_{n} = ar^{n - 1} \] so \[T_{6}: -\frac{1}{64} = \frac{1}{2}\times r^{6 - 1}\] solving for r is the task divide both sides by 1/2 \[-\frac{1}{32} = r^5..... or ...... r = \sqrt[5]{-\frac{1}{32}}....... r = - \frac{1}{2}\]

OpenStudy (campbell_st):

so the terms are \[\frac{1}{2}, \frac{1}{2} \times - \frac{1}{2}, - \frac{1}{4} \times - \frac{1}{2}.... \]